OpenStudy (sparklestaraa):

A scientist is studying the growth of a particular species of plant. He writes the following equation to show the height of the plant f(n), in cm, after n days:

OpenStudy (anonymous):

hummmm hey do u need help I am avaliable

OpenStudy (sparklestaraa):

f(n) = 12(1.03)^n @saynabahmed13 Yes yes i need help :D Part A: When the scientist concluded his study, the height of the plant was approximately 16.13 cm. What is a reasonable domain to plot the growth function? Part B: What does the y-intercept of the graph of the function f(n) represent? Part C: What is the average rate of change of the function f(n) from n = 3 to n = 10, and what does it represent?

OpenStudy (sparklestaraa):

@misty1212 I need help with this and the last part to another one :D

OpenStudy (misty1212):

hi!!

OpenStudy (misty1212):

did you get A, B?

OpenStudy (sparklestaraa):

no i didn't :(

OpenStudy (misty1212):

\[f(n) = 12(1.03)^n\] \[f(n) = 12(1.03)^n=16.13\] solve for \(n\) by checking

OpenStudy (sparklestaraa):

wait the full question isn't there

OpenStudy (sparklestaraa):

that needs to be here too

OpenStudy (misty1212):

looks like \(n=10\) check this http://www.wolframalpha.com/input/?i=12%281.03%29%5E10

OpenStudy (sparklestaraa):

ahh ok so is that for a b c d or e f

OpenStudy (misty1212):

wow lets go slow

OpenStudy (misty1212):

Part A: When the scientist concluded his study, the height of the plant was approximately 16.13 cm. What is a reasonable domain to plot the growth function?

OpenStudy (misty1212):

solve as above, you see \(n=10\) so ten weeks

OpenStudy (sparklestaraa):

ohhh so its the domain is 0 - 10 ?

OpenStudy (misty1212):

yes

OpenStudy (misty1212):

Part B: What does the y-intercept of the graph of the function f(n) represent?

OpenStudy (sparklestaraa):

wait is it any real number between 1- 10

OpenStudy (misty1212):

the y intercept is what you get when \(n=0\) which is \(f(0)=12(1.03)^0=12\)` \

OpenStudy (misty1212):

that is what you start with, in other words the plant was 12 cm tall to start

OpenStudy (misty1212):

Part C: What is the average rate of change of the function f(n) from n = 3 to n = 10, and what does it represent?

OpenStudy (misty1212):

\[\frac{f(10)-f(3)}{10-3}\] is the average rate of change lets use a caclulator

OpenStudy (sparklestaraa):

ok :)

OpenStudy (misty1212):

http://www.wolframalpha.com/input/?i=%2812%281.03%29%5E10-12%281.03%29%5E3%29%2F7 looks like about \(.43\)

OpenStudy (misty1212):

on average it grew \(.43\) cm per day from day 3 to day 10

OpenStudy (sparklestaraa):

why is it day 3 to 10? @misty1212

OpenStudy (sparklestaraa):

Misty what is a and e and f ???